For nos. 9-10, refer to the problem below:

Choc-o is sold in packets with an advertised mean weight of 0.5kgs. The standard deviation is known to

be 0.11 kilograms. A consumer group wishes to check the accuracy of the advertised mean and takes a

sample of 36 packets finding an average weight of 0.47kgs.

______ 9. What test statistic should be used?

A. population mean B. t-test C. z-test D. standard deviation

______ 10. What is the computed value?

A. – 1. 636 B. -1.488 C. 0. 833 D. 5. 551


1
Expert's answer
2021-05-03T07:06:00-0400

9.As the sample size is 36 So Generally z-test satistics is used.


10."=\dfrac{(x-\mu)\sqrt{n}}{\sigma}" "=\dfrac{(0.47-0.5)\sqrt{36}}{0.11}=\dfrac{-0.03\times 6}{0.11}=-1.636"

Hence The computed value is -1.636.